OpenStudy (anonymous):

Find the area under the curve given by the function f(x)=x^3+x^2-a on the interval [-2,1]

OpenStudy (anonymous):

are under the curve is found by integrating f(x) and evaluating the integral over the region between x = -2 and x=1 (In other words, evaluate the integral of f(x) at x=1 and subtract the value of the integral of f(x) at x = -2)

OpenStudy (anonymous):

Just tried to reference the text, but I'm having trouble with integrals

OpenStudy (anonymous):

I'm not an expert, but on one like this, set it up like this: \[ f(x)=x^3+x^2-a\] \[\int\limits_{-2}^{1} f(x)dx = \int\limits_{-2}^{1}(x^3+x^2-a)dx\]

OpenStudy (anonymous):

when you integrate a term like x^3, you end up with a term (x^4)/4 The way to see this is that if you differentiated (x^4)/4, you would get 4(x^3)/4 which simplifies to x^3. So integrating is finding the term that would to allow you to differentiate to get back to the original term. On relatively easy ones like this, a shortcut is to increase the exponent by 1 and then divide the new term by the value of the exponent.

OpenStudy (anonymous):

OK, I think I'm tracking...

OpenStudy (anonymous):

good... it takes me a bit to type all this :) so it's great if you can pick up the concept from what I've said so far... Do you want to post what you get? I will help check it :)

OpenStudy (anonymous):

@JakeV8 Nice explained :)

OpenStudy (anonymous):

** Nicely.

OpenStudy (anonymous):

yea, I will try... Thanks for all the help!

OpenStudy (anonymous):

It's going to take me a while to work through my references though...so please don't wait on my account though. Thanks for the steer in the right direction... :)